# Bifurcation mechanism of quasihalo orbit from Lissajous orbit

### v2

@@ Abstract
This paper presents a general analytical method to describe the center manifolds of collinear libration points in the Restricted Three-body Problem (RTBP). It is well-known that these center manifolds include Lissajous orbits, halo orbits, and quasihalo orbits. Previous studies have traditionally treated these orbits separately by iteratively constructing high-order series solutions using the Lindstedt-Poincaré method. Instead of relying on resonance between their frequencies, this study identifies that halo and quasihalo orbits arise due to intricate coupling interactions between in-plane and out-of-plane motions. To characterize this coupling effect, a novel concept, coupling coefficient $\eta$, is introduced in the RTBP, incorporating the coupling term $\eta\Delta x$ in the $z$-direction dynamics equation, where $\Delta$ represents a formal power series concerning the amplitudes. Subsequently, a uniform series solution for these orbits is constructed up to a specified order using the Lindstedt-Poincaré method. For any given paired in-plane and out-of-plane amplitudes, the coupling coefficient $\eta$ is determined by the bifurcation equation $\Delta=0$. When $\eta$ = 0, the proposed solution describes Lissajous orbits around libration points. As $\eta$ transitions from zero to non-zero values, the solution describes quasihalo orbits, which bifurcate from Lissajous orbits. Particularly, halo orbits bifurcate from planar Lyapunov orbits if the out-of-plane amplitude is zero. The proposed method provides a unified framework for understanding these intricate orbital behaviors in the RTBP.
@@ 1 Introduction
The Restricted Three-body Problem (RTBP) serves as a fundamental dynamical model to study the motion of asteroids or satellites under the gravitational influence of two primaries [1, 2, 3]. The RTBP has five equilibrium points called the Lagrange points $L_{i}$ ($i$ = 1, 2, …, 5), among which three so-called collinear libration points are unstable, and the remaining two are triangular libration points. The linear behavior of three collinear libration points is of the type center $\times$ center $\times$ saddle [4, 5, 6]. The center manifold of three collinear libration points includes Lissajous orbits around libration points, halo orbits, and quasihalo orbits around halo orbits [7]. These libration point orbits play a key role in astronomy and many space science missions [8, 9, 10, 11].
@@
Numerical methods (numerical integration, optimization technology) and semi-analytical methods (parameterization method, normal form, Lindstedt-Poincaré method) are often used to study the dynamics in the center manifolds of collinear libration points. In most cases, numerical methods still require accurate analytical initial values. On the other hand, high-order analytical solutions explicitly and accurately describe the dynamics around libration points, which is very useful for the preliminary space mission design. Farquhar initially proposed the third-order solution for small-amplitude Lissajous orbits and introduced the concept of “halo orbit” [12, 13]. Richardson [14] presented a third-order analytical solution of halo orbits around collinear libration points in the RTBP. He employed the Lindstedt-Poincaré method to construct this solution by introducing a correction term to modify the out-of-plane frequency. By means of the Lindstedt-Poincaré method, Jorba and Masdemont [4, 15] obtained high-order series solutions of Lissajous orbits and halo orbits around collinear libration points with the help of ad hoc algebraic manipulators. Celletti and Luo et al. [16, 17] conducted an analytical study of the Lissajous and halo orbits around collinear points in RTBP for arbitrary mass ratios, utilizing a canonical transformation procedure. Paez and Guzzo [18] provided a semi-analytical construction of halo orbits and halo tubes in the elliptic RTBP by implementing a nonlinear Floquet–Birkhoff resonant normal form. However, these analytical methods cannot yield high-order series solutions of quasihalo orbits. Gómez et al. [19] took a halo orbit with a constant amplitude as a backbone (periodic-varying equilibrium point), and further employed Floquet transformation and Lindstedt-Poincaré method to construct quasihalo orbits. This method necessitates specifying the magnitude of the reference halo orbit and using Floquet transformation, and does not directly provide a complete analytical expression for quasihalo orbits.
@@
So far, there is no explicit solution to uniformly describe Lissajous orbits, halo orbits and quasihalo orbits in the center manifolds, although some semi-analytical methods, such as normal form [4] and parameterization methods [6], or numerical methods, such as multiple Poincaré section [20], can provide a comprehensive description for the center manifolds of RTBP. References [14, 4, 15] successfully gave the third or higher-order series solutions of the Lissajous and halo orbits, yet they failed to construct a high-order series solution of the quasihalo orbits. Since the discovery of halo orbits in 1969, the astrodynamics community has traditionally attributed their origin to bifurcations from planar Lyapunov orbits when the in-plane and out-of-plane frequencies are in a 1:1 resonance. In this paper, we also contend that halo orbits indeed arise from bifurcations of planar Lyapunov orbits. However, we employ a new mechanism, coupling-induced bifurcation, to explain the generation of halo and quasihalo orbits. Consider a two-degree-of-freedom nonlinear dynamical system similar to the part of the center manifolds in the RTBP, but in this system, the degrees of freedom are decoupled, i.e., ${\left[{\ddot{x},\ddot{y}}\right]^{\rm{T}}}={\left[{f(x),f(y)}\right]^{\rm{T}}}$. It is evident that bifurcations of periodic orbits do not occur even under frequency resonance. Drawing a distinction between this hypothetical dynamical system and the RTBP, it can be inferred that the true trigger for the emergence of halo orbits through planar Lyapunov orbit bifurcation is the nonlinear coupling effect between the two degrees of freedom, rather than a resonance between the two frequencies. Considering the nonlinear coupling of in-plane and out-of-plane motions, this paper introduces the concept of a coupling coefficient $\eta$ into the RTBP for the first time. Additionally, a new correction term $\eta\Delta x$ is incorporated into the RTBP equation to construct a unified semi-analytical solution for the center manifolds of collinear libration points in the RTBP. Here, $\Delta$ is a power series concerning in-plane and out-of-plane amplitudes. When the coupling coefficient $\eta$ is zero, the correction item is inactive, and the series solution describes Lissajous orbits around libration points. Otherwise, the series solution describes quasihalo orbits, of course, including halo orbits. In this way, the center manifolds of collinear libration points in the RTBP are uniformly described with a high-order series solution.
@@
The contribution of this paper lies in the novel proposition of the coupling-induced bifurcation mechanism to explain the generation of quasihalo orbits. This marks the first complete approximate analytical solution for quasihalo orbits, providing a comprehensive analytical solution for the entire central manifold, including Lissajous orbits, halo orbits, and quasihalo orbits. Additionally, the practical convergence domain of the Lindstedt-Poincaré expansion of quasihalo orbits is presented. Lastly, the proposed method transforms the dynamic orbit bifurcation problem of vector fields into a static bifurcation problem for the solutions of the bifurcation equation. This method is applicable for general analytical bifurcation analysis in dynamical systems.
@@ 2 Dynamical model
This section introduces the dynamical model of the classical RTBP. This model serves as a good approximation for the motion of an infinitesimal particle (asteroids or spacecraft) under the gravitational attraction of two primaries. The attraction of the particle on the primaries is neglected so the two primaries rotating around their common center of mass in a Kepler orbit. Here, our focus lies on the circular orbit, i.e., a Kepler orbit with zero eccentricity. The motion of the particle is usually described in a synodic coordinate system. In this system, the origin is positioned at the centroid of the two primaries. The $X$-axis points from the smaller primary to larger primary, the $Z$-axis is perpendicular to the plane of the Kepler orbit and positive in the direction of the angular momentum, and the $Y$-axis completes a right-hand triad. Consequently, the position vector of the smaller primary and larger primary is ($-\mu$, 0, 0) and ($1-\mu$, 0, 0), respectively. Then, the differential equation governing the motion of the particle in the synodic coordinate system is expressed as follows [21, 7]:
@@
As mentioned in the Introduction section, the RTBP has five equilibrium points. In this paper, we focus on the three collinear libration points $L_{1}$, $L_{2}$, and $L_{3}$. Let $\gamma_{i}(i=1,2,3)$ denote the distance from $L_{i}$ to the closet primary. This distance is determined by the unique positive root of the Euler quantic equation [7],
@@
where $\rho^{2}=x^{2}+y^{2}+z^{2}$, $P_{n}$ is Legendre polynomials, and the constant coefficients $c_{n}(\mu)$ depend on the system parameters $\mu$,
@@
The left-hand linear terms form the foundational components of the series solution, while the right-hand nonlinear ones are crucial for constructing halo and quasihalo orbits. They are well-known as bifurcations from large-amplitude planar Lyapunov orbit.
@@ 3 Analytical construction of center manifolds in RTBP
In this section, the coupling coefficient is precisely defined. A coupling correction term is introduced to the original dynamical equation, which leads to a modified linear solution that accounts for coupling effects. Based on the new linear solution, the Lindstedt–Poincaré method is applied to construct a unified analytical solution for the center manifold near collinear libration points in the RTBP.
@@ 3.1 Form of the analytical solution
${\omega_{0}}=\sqrt{\frac{{2-{c_{2}}+\sqrt{9c_{2}^{2}-8{c_{2}}}}}{2}},{v_{0}}=\sqrt{{c_{2}}},\kappa=-\frac{{{\omega_{0}}^{2}+1+{2c_{2}}}}{{2{\omega_{0}}}}$ (13)
@@
where $d_{00}=c_{2}-{\omega_{0}}^{2}\neq 0$ is a constant correction factor and $\eta d_{00}=0$. It is evident that in the linear case, $\eta d_{00}=0$ only has a unique trivial solution, i.e., $\eta=0$, indicating the absence of coupling between the in-plane and out-of-plane motions. Halo/quasihalo orbits only appears when the coupling coefficient $\eta$ is not-zero in higher-order solutions. Therefore, we extend (15) to higher-order case by introducing a higher-order correction term to the third equation of (9) as follows:
@@
where the higher-order term $\eta\Delta x$ is the product of the coupling coefficient $\eta$, the nonlinear correction factor $\Delta$, and $x$, satisfying $\Delta=\sum\limits{{d_{ij}}{\alpha^{i}}{\beta^{j}}}$ and $\eta\Delta=0$. Since the coefficients $d_{ij}$ of $\Delta$ also include $\eta$, $\Delta$ is a polynomial function of $\eta$. During the process, non-trivial solutions of $\eta$ are determined from the condition $\Delta=0$. When the nonlinear terms in (16) are taken into account, the semi-analytical solution is formulated as formal expansions in powers of $\alpha$ and $\beta$,
@@
where ${\theta_{1}}{\rm{=}}{\omega}t+{\varphi_{1}}$,${\theta_{2}}{\rm{=}}\nu t+{\varphi_{2}}$. Considering the presence of high-order terms, the frequencies should also be expanded in the power series of $\alpha$ and $\beta$,
@@
which provides the implicit relationship between $\eta$ and $\alpha$ and $\beta$, i.e., $\eta=\eta(\alpha,\beta)$. Equations (17), (18), and (19) together constitute a comprehensive analytical description of the central manifolds around the collinear libration in the RTBP. The process for solving the coefficients $x_{ijkm}$, $y_{ijkm}$, $z_{ijkm}$, $\omega_{ij}$, $v_{ij}$ and $d_{ij}$ is provided in Appendix 3.2, and it is similar to the approach in [4], except for the parts related to the coupling coefficient $\eta$.
When $\eta=0$, the solution (17) represents Lissajous orbits. Specifically, if $\beta=0$, it corresponds to planar Lyapunov orbits; if $\alpha=0$, it corresponds to vertical Lyapunov orbits. When $\eta\neq 0$, the solution (17) represents quasihalo orbits. Particularly, if $\eta>0$ $(\eta<0)$ and $\beta=0$, it results in north (south) halo orbits. No solutions exist if $\alpha=0$, which will be demonstrated in the next section. In summary, the high-order series solution (17) uniformly describes the center manifolds of collinear libration points in the RTBP.
Although the analysis above only considered the coupling of $x$ applied to $z$, different forms of coupling will actually bifurcate into different families of orbits. When the $x$ coupling effect is applied to $z$, the solution of $\Delta=0$ indicates that this family of orbits bifurcates from planar Lyapunov orbits (and the corresponding Lissajous orbit), resulting in the so-called halo and quasihalo orbits. It is known that halo orbits cannot bifurcate from vertical Lyapunov orbits in the RTBP. Therefore, if we consider coupling $z$ to $x$, the equation $\Delta=0$ has no solution. Furthermore, if we consider coupling $\eta\Delta y$ to $z$ or $\eta\Delta z$ to $y$, another family of orbits, known as axis orbit and quasi-axis orbits, will be obtained. It is well-known that axis orbits can be regarded as bifurcating either from planar Lyapunov orbits or from vertical Lyapunov orbits [23].
@@ 3.2 Solving for undetermined coefficients
This part provides the process for solving the coefficients $x_{ijkm}$, $y_{ijkm}$, $z_{ijkm}$, $\omega_{ij}$, $v_{ij}$ and $d_{ij}$. From (17) and (18) and (19), it is known that obtaining the analytical expressions for the center manifold near the $L_{1}$ point is contingent upon determining the coefficients $x_{ijkm}$, $y_{ijkm}$, $z_{ijkm}$, $\omega_{ij}$, $v_{ij}$ and $d_{ij}$. Utilizing the Lindstedt–Poincaré method with iterations starting from the linear solution allows for the computation of these coefficients. Now, if the coefficients and the corresponding solution up to order $n-1$ are established, substituting this solution into the right-hand side of (16) yields the three known series up to order $n$, denoted by $p$, $q$, and $r$. The terms up to the $n$-order from these series, $p_{ijkm}$, $q_{ijkm}$, and $r_{ijkm}$ ($i+j=n$), are selected for solving $n$-order unknown coefficients.
Then, the known and unknown terms on the left side of the equation are classified. The first and second derivatives of x are expressed as:
Analogous operations can be carried out for $y$ and $z$. For the first-order derivatives and the coupling correction term $\Delta x$, their known and unknown terms can be classified as shown in Table 1, where $i_{f}$ represent the power of frequency terms or coupling term $\Delta$, and $i_{g}$ represents the power of coordinate terms, $\delta_{ij}$ denotes Kronecker function. A similar classification for the second-order derivatives is provided in Table 2.
Table 1: Classification for the derivatives of $x$ and $y$ and coupling correction term $\Delta x$.
@@
Table 2: Classification for the second derivatives of $x$, $y$, and $z$.
@@
Then, all the known terms from Table 1 and Table 2 are combined into the known series on the right-hand side of (16) and they are denoted as ${\bar{p}_{ijkm}}$, ${\bar{q}_{ijkm}}$ and ${\bar{r}_{ijkm}}$. Additionally, for the unknown terms $\Omega_{i-1j}$ and ${\rm N}_{ij-1}$, they also include some known terms, i.e., $\Omega_{i-1j}=2\omega_{00}\omega_{i-1j}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}$, ${\rm N}_{ij-1}=2v_{00}v_{ij-1}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{v_{{i_{1}}{j_{1}}}}{v_{{i_{2}}{j_{2}}}}}$, where $i_{1}+j_{1}+i_{2}+j_{2}=n-1$ and $C_{n}^{k}$ denotes a combination. These known terms need to be merged into the right-hand side of the equation as well. In conclusion, the linear system of equations for each term of order $n$ is as follows:
$\begin{array}[]{l}-(\varpi_{km}^{2}+1+2{c_{2}}){x_{ijkm}}-2{\varpi_{km}}{y_{xjkm}}-2\left({{\omega_{00}}+\kappa}\right){\omega_{i-1j}}{\delta_{1k}}{\delta_{0m}}={{\bar{\bar{p}}}_{ijkm}}\\ -2{\varpi_{km}}{x_{ijkm}}+({c_{2}}-1-\varpi_{km}^{2}){y_{xjkm}}-2\left({\kappa{\omega_{00}}+1}\right){\omega_{i-1j}}{\delta_{1k}}{\delta_{0m}}={{\bar{\bar{q}}}_{ijkm}}\\ ({c_{2}}-\varpi_{km}^{2}){z_{ijkm}}-{d_{00}\eta}{x_{ijkm}}-2{v_{00}}{v_{ij-1}}{\delta_{0k}}{\delta_{1m}}-2\eta{\omega_{00}}{\omega_{i-1j}}{\delta_{1k}}{\delta_{0m}}-\eta{d_{i-1j}}{\delta_{1k}}{\delta_{0m}}={{\bar{\bar{r}}}_{ijkm}}\end{array}$ (21)
where ${\varpi_{km}}=k{\omega_{00}}+m{v_{00}}$, and

### v3

@@ Abstract
This paper presents a general analytical method to describe the center manifolds of collinear libration points in the Restricted Three-body Problem (RTBP). It is well-known that these center manifolds include Lissajous orbits, halo orbits, and quasihalo orbits. Previous studies have traditionally tackled these orbits separately by iteratively constructing high-order series solutions using the Lindstedt-Poincaré method. Instead of relying on resonance between their frequencies, this study identifies that halo and quasihalo orbits arise due to intricate coupling interactions between in-plane and out-of-plane motions. To characterize this coupling effect, a novel concept, coupling coefficient $\eta$, is introduced in the RTBP, incorporating the coupling term $\eta\Delta x$ in the $z$-direction dynamics equation, where $\Delta$ represents a formal power series concerning the amplitudes. Subsequently, a uniform series solution for these orbits is constructed up to a specified order using the Lindstedt-Poincaré method. For any given paired in-plane and out-of-plane amplitudes, the coupling coefficient $\eta$ is determined by the bifurcation equation $\Delta=0$. When $\eta$ = 0, the proposed solution describes Lissajous orbits around libration points. As $\eta$ transitions from zero to non-zero values, the solution describes quasihalo orbits, which bifurcate from Lissajous orbits. Particularly, halo orbits bifurcate from planar Lyapunov orbits if the out-of-plane amplitude is zero. The proposed method provides a unified framework for understanding these intricate orbital behaviors in the RTBP.
keywords:
Restricted three-body problem , Center manifold , Lissajous orbit , Quasihalo orbit , Coupling coefficient
journal: Arvix
[…]
[inst1]organization=Advanced Institute for Materials Research, Tohoku University,addressline=2-1-1 Katahira, Aoba-ku, city=Sendai, postcode=980-8577, country=Japan
@@ 1 Introduction
The Restricted Three-body Problem (RTBP) serves as a fundamental dynamical model to study the motion of asteroids or satellites under the gravitational influence of two primaries [1, 2, 3]. The RTBP has five equilibrium points called the Lagrange points $L_{i}$ ($i$ = 1, 2, …, 5), among which three so-called collinear libration points are unstable, and the remaining two are triangular libration points. The linear behavior of three collinear libration points is of the type center × center × saddle. The center manifold of three collinear libration points includes Lissajous orbits around libration points, halo orbits, and quasihalo orbits around halo orbits [4]. These libration point orbits play a key role in astronomy and many space science missions [5, 6, 7, 8].
@@
Numerical methods (numerical integration, optimization technology) and analytical methods (parameterization method, norm form, Lindstedt-Poincaré method) are often used to study the dynamics in the center manifolds of collinear libration points. In most cases, numerical methods still require accurate analytical initial values. On the other hand, high-order analytical solutions explicitly and accurately describe the dynamics around libration points, which is very useful for the preliminary space mission design. Farquhar initially proposed the third-order solution for small-amplitude Lissajous orbits and introduced the concept of “halo orbit” [9, 10]. Richardson [11] presented a third-order analytical solution of halo orbits around collinear libration points in the RTBP. He employed the Lindstedt-Poincaré method to construct this solution by introducing a correction term to modify the out-of-plane frequency. By means of the Lindstedt-Poincaré method, Jorba and Masdemont [12, 13] obtained high-order series solutions of Lissajous orbits and halo orbits around collinear libration points with the help of ad hoc algebraic manipulators. Celletti and Luo et al. [14, 15] conducted an analytical study of the Lissajous and halo orbits around collinear points in RTBP for arbitrary mass ratios, utilizing a canonical transformation procedure. Paez and Guzzo [16] provided a semi-analytical construction of halo orbits and halo tubes in the elliptic RTBP by implementing a nonlinear Floquet–Birkhoff resonant normal form. However, these analytical methods cannot yield high-order series solutions of quasihalo orbits. Gómez et al. [17] took a halo orbit with a constant amplitude as a backbone (periodic-varying equilibrium point), and further employed Floquet transformation and Lindstedt-Poincaré method to construct quasihalo orbits. This method necessitates specifying the magnitude of the reference halo orbit and using Floquet transformation, and does not directly provide a complete analytical expression for quasihalo orbits.”.
@@
So far, there is no semi-analytical method to uniformly describe Lissajous orbits, halo orbits and quasihalo orbits in the center manifolds, although some numerical methods, such as norm form [12] and multiple Poincaré section [18], can provide a comprehensive description for the center manifolds of RTBP. References [11, 12, 13] successfully gave the third or higher-order series solutions of the Lissajous and halo orbits, yet they failed to construct a high-order series solution of the quasihalo orbits. Since the discovery of halo orbits in 1969, the astrodynamics community has traditionally attributed their origin to bifurcations from planar Lyapunov orbits when the in-plane and out-of-plane frequencies are in a 1:1 resonance. In this paper, we also contend that halo orbits indeed arise from bifurcations of planar Lyapunov orbits. However, We employ a new mechanism, coupling-induced bifurcation, to explain the generation of halo and quasihalo orbits. Consider a two-degree-of-freedom nonlinear dynamical system similar to the part of the center manifolds in the RTBP, but in this system, the degrees of freedom are decoupled, i.e., ${\left[{\ddot{x},\ddot{y}}\right]^{\rm{T}}}={\left[{f(x),f(y)}\right]^{\rm{T}}}$. It is evident that bifurcations of periodic orbits do not occur even under frequency resonance. Drawing a distinction between this hypothetical dynamical system and the RTBP, it can be inferred that the true trigger for the emergence of halo orbits through planar Lyapunov orbit bifurcation is the nonlinear coupling effect between the two degrees of freedom, rather than a resonance between the two frequencies. Considering the nonlinear coupling of in-plane and out-of-plane motions, this paper introduce the concept of a coupling coefficient $\eta$ into the RTBP for the first time. Additionally, a new correction term $\eta\Delta x$ is incorporated into the RTBP equation to construct a unified semi-analytical solution for the center manifolds of collinear libration points in the RTBP. Here, $\Delta$ is a power series concerning in-plane and out-of-plane amplitudes. When the coupling coefficient $\eta$ is zero, the correction item is inactive, and the series solution describes Lissajous orbits around libration points. Otherwise, the series solution describes quasihalo orbits, of course, including halo orbits. In this way, the center manifolds of collinear libration points in the RTBP are uniformly described with a high-order series solution.
@@
The contribution of this paper lies in the novel proposition of the coupling-induced bifurcation mechanism to explain the generation of quasihalo orbits. This marks the first complete approximate analytical solution for quasihalo orbits, providing a comprehensive analytical solution for the entire central manifold, including Lissajous orbits, halo orbits, and quasihalo orbits. Additionally, the practical convergence domain of quasihalo orbits is presented for the first time. Lastly, the proposed method transforms the dynamic orbit bifurcation problem of vector fields into a static bifurcation problem for the solutions of the bifurcation equation. This method is applicable for general analytical bifurcation analysis in dynamical systems.
@@ 2 Dynamical model
This section introduces the dynamical model of the classical RTBP. This model serves as a good approximation for the motion of an infinitesimal particle (asteroids or spacecraft) under the gravitational attraction of two primaries. The attraction of the particle on the primaries is neglected so the two primaries rotating around their common center of mass in a Kepler orbit. Here, our focus lies on the circular orbit, i.e., a Kepler orbit with zero eccentricity. The motion of the particle is usually described in a synodic coordinate system. In this system, the origin is positioned at the centroid of the two primaries. The $X$-axis points from the smaller primary to larger primary, the $Z$-axis is perpendicular to the plane of the Kepler orbit and positive in the direction of the angular momentum, and the $Y$-axis completes a right-hand triad. Consequently, the position vector of the smaller primary and larger primary is ($-\mu$, 0, 0) and ($1-\mu$, 0, 0), respectively. Then, the differential equation governing the motion of the particle in the synodic coordinate system is expressed as follows [19]:
@@
As mentioned in the Introduction section, the RTBP has five equilibrium points. In this paper, we focus on the three collinear libration points $L_{1}$, $L_{2}$, and $L_{3}$. Let $\gamma_{i}(i=1,2,3)$ denote the distance from Li to the closet primary. This distance is determined by the unique positive root of the Euler quantic equation [4, 19],
@@
where $\rho=x^{2}+y^{2}+z^{2}$, $P_{n}$ is Legendre polynomials, and the constant coefficients $c_{n}(\mu)$ are depend on the system parameters $\mu$,
@@
The left-hand linear terms form the foundational components of the series solution, while the right-hand nonlinear ones are crucial for constructing halo and quasihalo orbits. They are well-known as bifurcations from large-amplitude planner Lyapunov orbit.
@@ 3 Analytical construction of center manifolds in RTBP
In this section, a semi-analytical solution for the center manifolds of collinear libration points in the RTBP is constructed using the Lindstedt–Poincaré method. Lindstedt–Poincaré method is an iterative computational technology that begins from the fundamental first-order solution of the system. Subsequently, it continuously adjusts the relationship between the frequency and amplitude to obtain a higher-order series solution by iterating the known low-order solution step by step.
@@ 3.1 Form of the analytical solution
${\omega_{0}}=\sqrt{\frac{{2-{c_{2}}+\sqrt{9c_{2}^{2}-8{c_{2}}}}}{2}},{v_{0}}=\sqrt{{c_{2}}},\kappa=-\frac{{{\omega_{0}}^{2}+1+{c_{2}}}}{{2{\omega_{0}}}}$ (13)
@@
where $d_{00}=c_{2}-{\omega_{0}}^{2}\neq 0$ is a constant correction factor and $\eta d_{00}=0$. It is evident that for linear equation (15), the only trivial solution is $\eta=0$, indicating the absence of coupling between the in-plane and out-of-plane motions. Halo/quasihalo orbits only appears when the coupling coefficient $\eta$ is not-zero in higher-order solutions. Therefore, we extend (15) to higher-order case by introducing a higher-order correction term to the third equation of (9) as follows:
@@
where the higher-order term $\eta\Delta x$ is the product of the coupling coefficient $\eta$, the nonlinear correction factor $\Delta$, and $x$, satisfying $\Delta=\sum\limits{{d_{ij}}{\alpha^{i}}{\beta^{j}}}$ and $\eta\Delta=0$. During the process, non-trivial solutions of $\eta$ are determined from the condition $\Delta=0$. When the nonlinear terms in (16) are taken into account, the semi-analytical solution is formulated as formal expansions in powers of $\alpha$ and $\beta$,
@@
where ${\theta_{1}}{\rm{=}}{\omega}t+{\varphi_{1}}$,${\theta_{2}}{\rm{=}}\nu t+{\varphi_{2}}$. Considering the nonlinear terms, the frequencies should also be expanded in the power series of $\alpha$ and $\beta$,
@@
which provides the implicit relationship between $\eta$ and $\alpha$ and $\beta$, i.e., $\eta=\eta(\alpha,\beta)$. Equations (17), (18), and (19) together constitute a comprehensive analytical description of the central manifolds around the collinear libration in the RTBP.
Remark 1. In fact, the above correction is not unique, as long as the correction term can represent the coupling between in-plane and out-of-plane motions, such as adding $\eta\Delta y$ to the third equation or $\eta\Delta z$ to the first (second) equation of (16).
Remark 2. When $\eta=0$, the solution (17) represents Lissajous orbits. Specifically, if $\beta=0$, it corresponds to planar Lyapunov orbits; if $\alpha=0$, it corresponds to vertical Lyapunov orbits. When $\eta\neq 0$, the solution (17) represents quasihalo orbit. Particularly, if $\eta>0$ $(\eta<0)$ and $\beta=0$, it results in north (south) halo orbits. No solutions exist if $\alpha=0$, which will be demonstrated in the next section. In sum up, the high-order series solution (17) uniformly describes the center manifolds of collinear libration points in the RTBP.
Remark 3. In (17) and (18), $i$ and $j\in\mathbb{N}$, $k$ and $m\in\mathbb{Z}$. Due to the symmetry of the RTBP, $x(t)$ and $z(t)$ are formulated as a cosine series, and $y(t)$ as a sine series. $p$ and $q$ have the same parity as $i$ and $j$, respectively. Besides, due to the symmetry of sine and cosine functions, it can be assumed that $m\geq 0$, and $n\geq 0$ when $m=0$. The series of $\omega$ and $v$ only include even items. These properties are useful for conserving computational storage and time.
@@ 3.2 Solving for undetermined coefficients
Now, our goal is to compute the coefficients $x_{ijkm}$, $y_{ijkm}$, $z_{ijkm}$, $\omega_{ij}$, $v_{ij}$ in (17) and (18) up to a finite order $n$. The Lindstedt-Poincaré method is utilized to calculate these coefficients following an iterative scheme from the linear solution. Compared to the solution of the linear part (14), we can determine $x_{1010}=1$, $y_{1010}=\kappa$, $z_{1010}=\eta$, $z_{0101}=1$, $\omega_{00}=\omega_{0}$, $v_{00}=v_{0}$. By substituting this linear solution into (16), the coefficients of the second-order solution can be derived. Similarly, when the coefficients up to order $n-1$ are obtained, i.e., $x(t)$, $y(t)$, and $z(t)$ are determined up to order $n-1$, $\omega$ and $v$ are determined up to order $n-2$. Substituting them into the right side of (16), we can obtain three power series up to order $n$, denoted by $p$, $q$, and $r$. Here, what we are interested in are those $n$-order terms. Without losing generality, the $n$-order terms of $p$, $q$, and $r$ are denoted by $p_{ijkm}$, $q_{ijkm}$, and $r_{ijkm}$ ($i+j=n$) respectively.
Table 1: Derivatives of $x$ and $y$ with respect to time and $\Delta x$.
@@
Table 2: Second derivatives of $x$, $y$, and $z$ with respect to time.
@@
Next, we analyze the composition of $n$-order terms on the left side of (16). According to (17), the derivatives of variable $x$ can be expressed as
Similarly, we can obtain the derivatives of $y$ and $z$. Let $fg$ denote the first-order derivative term, where $f$ represents the power series of the frequencies ($\omega$ and $v$), and $g$ represents the coordinate variables ($x$, $y$, and $z$). $(ij)_{f}$ and $(ij)_{g}$ denote their corresponding order. Then, the $fg$ satisfying $(ij)_{f}+(ij)_{g}=n$ constitutes the $n$-order terms we require. When $(ij)_{f}$ is 0 or $n-1$, the corresponding $(ij)_{g}$ is $n$ or 1 and $fg$ is an unknown term that needs to be solved. When $(ij)_{f}=1,2,...,n-2$, $fg$ is a known term that needs to be moved to the right-hand side of (16). Table 1 summarizes the unknown and known terms of the first derivatives and the product of $\Delta$ and $x$, where $\delta_{ij}$ denotes the Kronecker function. For the second derivatives of $x$, $y$, and $z$, it can be similarly summarized as shown in Table 2.
Then we move all the known terms to the right-hand side of (16), add them to $p_{ijkm}$, $q_{ijkm}$, and $r_{ijkm}$, and re-denote them as ${\bar{p}_{ijkm}}$, ${\bar{q}_{ijkm}}$ and ${\bar{r}_{ijkm}}$. Besides, we need to address the calculation of the unknown term $\Omega_{i-1j}$ and ${\rm N}_{ij-1}$. In fact, they are composed of the unknown term $2\omega_{00}\omega_{i-1j}$ ($2v_{00}v_{ij-1}$) and the remaining known terms, i.e., $\Omega_{i-1j}=2\omega_{00}\omega_{i-1j}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}$, ${\rm N}_{ij-1}=2v_{00}v_{ij-1}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{v_{{i_{1}}{j_{1}}}}{v_{{i_{2}}{j_{2}}}}}$, where $i_{1}+j_{1}+i_{2}+j_{2}=n-1$ and $C_{n}^{k}$ denotes a combination. Similar results can be obtained for the second derivatives of $y$ and $z$. In summary, the linear equation of n-order unknown coefficients is yielded
$\begin{array}[]{l}-(\varpi_{km}^{2}+1+2{c_{2}}){x_{ijkm}}-2{\varpi_{km}}{y_{xjkm}}-2\left({{\omega_{00}}+\kappa}\right){\omega_{i-1j}}{\delta_{1k}}{\delta_{0m}}={{\bar{\bar{p}}}_{ijkm}}\\ -2{\varpi_{km}}{x_{ijkm}}+({c_{2}}-1-\varpi_{km}^{2}){y_{xjkm}}-2\left({\kappa{\omega_{00}}+1}\right){\omega_{i-1j}}{\delta_{1k}}{\delta_{0m}}={{\bar{\bar{q}}}_{ijkm}}\\ ({c_{2}}-\varpi_{km}^{2}){z_{ijkm}}-{d_{00}\eta}{x_{ijkm}}-2{v_{00}}{v_{ij-1}}{\delta_{0k}}{\delta_{1m}}-2\eta{\omega_{00}}{\omega_{i-1j}}{\delta_{1k}}{\delta_{0m}}-{d_{i-1j}}{\delta_{1k}}{\delta_{0m}}={{\bar{\bar{r}}}_{ijkm}}\end{array}$ (21)
where ${\varpi_{km}}=k{\omega_{00}}+m{v_{00}}$, ${d_{00}}=\left({{c_{2}}-\omega_{00}^{2}}\right)\eta$, and